OpenStudy (anonymous):

can someone explain how to do this? Which of the following ordered pairs is the solution to the system of linear equations? x + y = 14 2x + y = 24

OpenStudy (anonymous):

-(x+y)=14*-1 -x-y=-14 + 2x+y=24 -------------- x=10 then plug in 10 for x and solve for y so you do: 10+y=14 y=14-10=4

OpenStudy (anonymous):

(10,4)

OpenStudy (anonymous):

okay thank you(: